OpenStudy (narissa):

evaluate 2(3x-y) when a)x=10 y=-3 b)x=-3 y=10 c)x=-3 y=-10

OpenStudy (narissa):

@Austin1617

OpenStudy (phi):

first do 3*x when x=10 what is 3*x ?

OpenStudy (narissa):

should the first one look like this 2(3+10+3)=66

OpenStudy (narissa):

30

OpenStudy (phi):

so now you have 2*(30-y) now do 30 - y when y is -3 30 - (-3) is what ?

OpenStudy (narissa):

is it -33

OpenStudy (phi):

two minuses make a plus 30 - -3 is the same as 30+3

OpenStudy (narissa):

oh 33*2 right?

OpenStudy (phi):

yes

OpenStudy (narissa):

66

OpenStudy (narissa):

thank u

OpenStudy (phi):

yes. do you get the general idea? for the next problem b)x=-3 y=10 we can swap in those numbers into 2(3x-y) to get 2*(3*-3 - 10)

OpenStudy (phi):

I put in the * to show multiplication because we know 3x means 3*x but if we put in -3 for x without the multiply sign: 3-3, that is wrong we really would want 3* -3 which is -9

OpenStudy (narissa):

-38?

OpenStudy (phi):

yes

OpenStudy (narissa):

thanks

OpenStudy (narissa):

would c look like this 2(-9+10=2

OpenStudy (phi):

yes, for c) you get 2*1= 2

OpenStudy (narissa):

thanks